# Evolution of boundedly rational learning in games

### v1

@@ Introduction
To explore how the players’ selection strengths $\beta$ might change over time, we use adaptive dynamics [38, 39, 40]. That is, we consider a large and well-mixed population of players (Fig. 1D). Each player randomly meets other population members to interact in a series of pairwise games. As a result, they obtain a payoff as described in the previous section on introspection dynamics. We define a player’s fitness to be its average payoff across all its interaction partners. Players with larger fitness are more likely to reproduce. The offspring of a reproducing player inherits the parent’s trait $\beta$ with probability $1\!-\!\mu$. With the complementary probability $\mu$, the offspring’s selection strength is slightly mutated, by adding some small noise to the parent trait. Overall, these assumptions define a dynamical process on the population level, with $\beta$ being the evolving trait. This process is straightforward to simulate. We give a detailed description of the algorithm in the Material and Methods, and provide the respective code online.
Figure 3: Adaptive dynamics of selection strength for three different social dilemmas. We explore the adaptive dynamics for three different social dilemmas: the prisoner’s dilemma, the snowdrift game, and the stag-hunt game. For each case, we depict a realization of an individual-based simulation (upper panels) and a pairwise invasibility plot (lower panels). Each simulation is initialized at a monomorphic population with a selection strength of $\beta\!=\!0$. Over time, mutations introduce additional selection strength values. Individuals obtain payoffs by randomly interacting with other population members, and learning strategies through introspection dynamics. They reproduce according to their payoff. For details, see Material and Methods. Pairwise invasibility plots illustrate the dynamics assuming a monomorphic resident population. They display (in color) which rare mutant traits ($y$-axis) have a positive invasion fitness $s_{x}(y)$, given the current resident ($x$-axis). Colored regions above the diagonal indicate that mutants with a higher trait value than the resident are favored to invade. Colored regions below the diagonal indicate evolution towards smaller trait values. Dashed lines indicate the position of singular trait values. A,B, For the depicted prisoner’s dilemma, we observe a dynamics of ever-increasing selection strengths. C,D, In the snowdrift game, the evolving $\beta$ values converge to a finite value. E,F, In the stag-hunt game, we observe evolutionary branching. After this occurs, there are two subpopulations. Members of the first subpopulation have low $\beta$ values, and hence they choose strategies mostly at random. Members of the second subpopulation exhibit high $\beta$ values; they tend to best respond to their respective opponent.
@@
Initially, we assume that the population is monomorphic, with a selection strength of $\beta\!=\!0$. To gain some intuition of the further dynamics, we revisit the previous two examples. Fig. 3A shows simulations for the prisoner’s dilemma. Here, we observe that $\beta$ evolves towards ever-increasing values. This outcome is consistent with the naive expectation that more effective learners should have a fitness advantage, and consistent with our analysis in Fig. 2A–D. For the other example, the snowdrift game, simulations are shown in Fig. 3C. Here, even though larger selection strengths are initially favored, the dynamics stops at some value well below $\beta\!=\!6$. That is, in contrast to our naive expectation – but consistent with our analysis in Fig. 2E–H – we observe the evolution of boundedly rational learning.
@@
In this formula, $\Pi_{x}(y)$ is the payoff of a mutant with selection strength $\beta\!=\!y$ in a resident population with $\beta\!=\!x$. This payoff is given by (3). The mutant can selectively invade the resident if and only if this invasion fitness is positive. In pairwise invasibility plots, the respective region for which $s_{x}(y)\!>\!0$ is highlighted in color. For the prisoner’s dilemma, this colored area comprises the entire space above the main diagonal (Fig. 3B). This means that independent of the exact selection strength $x$ of the resident, a mutant with trait $y$ can invade if and only if $y\!>\!x$. This implies the evolution of ever-increasing selection strengths.
@@
In contrast, in the snowdrift game in Fig. 3D, mutants with a higher selection strength are only favored when the resident’s selection strength is sufficiently small (if $x\!<\!x^{*}$ for some threshold $x^{*}$). We can numerically compute this threshold by setting the so-called selection gradient equal to zero (see Material and Methods). For the depicted example, we obtain $x^{*}\!=\!5.58$, in line with the evolutionary simulations. The pairwise invasibility plot also indicates that this singular point is evolutionarily stable. Once the population settles at a selection strength of $\beta\!=\!x^{*}$, neither a more effective nor a less effective learner can invade (as indicated by the white neighborhood around the vertical dashed line).
@@
Interestingly, these two examples of an ever-increasing selection strength and an evolutionarily stable finite selection strength are not the only possibilities. Instead, Fig. 3E shows a third scenario, using a stag-hunt game. Here, we observe evolutionary branching. At first, the monomorphic population again approaches a singular point, in this case at $x^{*}\!=\!3.01$. Once there, both mutant types can invade, those that are more effective learners and those that are less effective learners (indicated by the colored neighborhood of the vertical dashed line in Fig. 3F). This is the fingerprint of a branching point. From here, the population starts diverging in opposite directions. There is a sub-population for which the selection strength increases, and another one for which it decreases. The subsequent dynamics is more complex to describe analytically; we do so in the Supporting Information and Fig. S2. In the main text, we provide some intuition for why branching occurs in the first place. To this end, it is again instructive to look at the stationary distribution of introspection dynamics, for two different selection strengths $\beta$ (see Fig. S3A–D). When both players use the same $\beta$, a larger $\beta$ makes players choose the less efficient DD equilibrium more often (because defection is risk-dominant). If one player instead has a low $\beta$, both players occasionally coordinate on the more efficient CC equilibrium, while often avoiding harmful mis-coordination on the CD-outcome.
@@
Once selection strength is allowed to evolve, its dynamics also has important implications on the population’s average payoff. For example, in the prisoner’s dilemma, as individuals become more effective in their learning process, they increasingly learn to adopt the individually optimal strategy of defection. As a result, overall payoffs decrease in time (Fig. S4, left column). For the other two examples, the long-run effects on payoffs are more advantageous. Both for the snowdrift and the stag-hunt game, we observe an initially negative but eventually positive trend in the players’ payoffs (Fig. S4, middle and right column). In particular, the evolutionary branching in the stag-hunt game turns out to be both individually and collectively beneficial.
@@
So far we have studied the evolution of selection strength for a few isolated $2\!\times\!2$ games. To identify more general patterns, in a next step we systematically vary the parameters $A$ and $B$ of the stage game. For each resulting game, we compute the numerically exact position of the singular point (if it exists), and its stability properties. The result is Fig. 4. Here, white regions correspond to cases without a singular point. For the corresponding stage games, we predict selection strengths to increase indefinitely. Colored regions correspond to cases with a finite singular point. Here, we distinguish orange regions (where the point is evolutionarily stable) and purple regions (where we observe branching). Based on this figure, we make three observations.
@@
Figure 4: Evolutionary dynamics across all social dilemmas. So far, we described three special instances of games; here we systematically repeat this analysis for all $2\!\times\!2$ games. For each pair of game parameters $A$ and $B$, we numerically check whether there exists a finite singular trait value $\beta^{*}$. If it exists, we explore whether the respective trait value is evolutionarily stable (orange) or a branching point (purple). The color gradient represents the position of the singular trait: the lighter, the larger the value of $\beta^{*}$. Regions without a finite singular trait are left white. Black symbols indicate the position of the three examples displayed in Fig. 3.
@@
First, even though most prisoner’s dilemmas lead to ever-increasing selection strengths (as in Fig. 3A), there are exceptions. These exceptions occur if the absolute value $|A|$ is small (if there is only a small temptation to defect when the co-player cooperates). Here, we observe evolution towards a finite value of $\beta$. The underlying mechanism is different from the mechanism in the snowdrift game discussed earlier. Because $|A|$ is small compared to $B$, a co-player is more likely to occasionally switch to cooperation if the focal player cooperates. This positive influence on the co-player’s cooperation rate can make it worthwhile to have a finite $\beta$, and to cooperate occasionally (Fig. S3E–H).
@@
Second, in the snowdrift game, we observe evolution towards a finite selection strength for all games with $B\!>\!A$. In addition, as we consider a wider range of parameter values than considered in Fig. 4, we find another region exhibiting evolution towards finite $\beta$, see Fig. S5. This second region is given by $B\!<\!A\!-\!1$; we provide a derivation of these conditions in the Supporting Information. In both regions, the mechanism favoring finite selection strengths is the same: By having lower selection strengths, learners ensure they spend more time in their individually preferred equilibrium.
@@
Finally, in the stag-hunt game, we only observe evolutionary branching when it is risk-dominant to defect (that is, for $A\!<\!B$). If instead cooperation is risk-dominant, the prevalence of CC increases as a function of the players’ selection strength. Mutual cooperation is the optimal outcome for both players; hence, an ever-increasing selection strength is individually and mutually beneficial.
@@
Above, we documented several cases with an unexpected dynamics (e.g., evolution towards finite selection strengths or evolutionary branching). In each case, the dynamics occurred because at least one player gained some strategic advantage from being a less effective learner. In general, however, one might expect that these strategic advantages are strongest in pairwise games. Among two players, one player’s (possibly irrational) behavior tends to have a comparably strong effect on how the co-player might react. This effect could wane in groups of size $N\!>\!2$. To further explore the robustness of our results, in the Supporting Information we study the dynamics of three classes of public goods games. The three classes differ in the shape of the public good’s benefit function. Benefits are either linear, they show diminishing returns [49], or they take the form of a threshold function [50, 51]. While the computational steps involved in analyzing the adaptive dynamics are the same as before, the mathematical expressions now become more intricate. Importantly, however, the qualitative dynamics are similar to before. Depending on the payoff function, again we recover the three previously observed cases of indefinite evolution, convergence to a finite evolutionarily stable state, or evolutionary branching (Fig. S6, for $N\!=\!3$).
@@ Discussion
Evolutionary game theory explores how strategies and traits change over time [1, 2, 3]. Such an evolutionary approach can be motivated in two ways. On the one hand, the respective dynamical equations may describe long-run biological evolution. Here, individuals with better strategies are assumed to reproduce more often [4, 5]. On the other hand, evolutionary models can also be interpreted as a representation of an individual’s short-run learning process. Such models assume that better strategies are more readily adopted [6, 7, 8, 9, 10, 11, 12]. Herein, we blend these two interpretations. We ask how a long-term biological process may shape the effectivity of individual learning. To this end, we study the evolution of a key quantity in stochastic learning models, the players’ selection strength. This quantity determines to which extent payoff advantages sway individuals to change their strategies. When selection strength is small, individuals tend to ignore their strategies’ payoffs [18, 19, 20, 21, 22]. Here, strategy choices are mostly determined by chance. As the selection strength becomes larger, individuals become increasingly sensitive to (even minor) payoff differences.
@@
Motivated by these observations, herein we interpret selection strength as an individual trait that measures how rational individuals act when updating their strategies. We use this perspective to study the evolution of boundedly rational learning in games. To this end, we consider a model with two time scales. In the short run, individuals have fixed but possibly different selection strengths. We explore the resulting learning processes with introspection dynamics [31, 32, 33, 34, 35, 36, 37]; for the $2\times 2$ games we focus on, this process is equivalent to the logit-response dynamics studied in economics [32, 52]. We use this short-term perspective to ask whether individuals with larger selection strengths are able to take advantage of their opponents.
@@
In the long run, we also allow the individuals’ selection strength to evolve. Here, naively one might expect ever-increasing selection strengths. After all, large selection strengths make an individual’s learning process more effective, allowing them to better distinguish profitable strategies from unprofitable ones. Surprisingly, however, we identify several important game classes in which selection strength converges towards a finite value. In some cases, ineffective learning seems to act as a commitment device. By learning more erratically, individuals commit themselves to use strategies that run counter to their immediate interests. In such a case, more effectively learning opponents may be swayed to react, and to coordinate on an equilibrium that eventually benefits the more ineffective learner. These findings have interesting parallels in the biology literature on the red-king effect [53, 54, 55]. That literature suggests that in mutualisms among two co-evolving species too, it is the slower evolving species that eventually has an evolutionary advantage.
More generally, our study contributes to a research area that explores the origins and consequences of boundedly rational behavior. Gigerenzer and Selten point out that there are two complementary directions to the study of bounded rationality, “psychological plausibility” and “ecological rationality” [56]. The first direction focuses on analyzing the mind (such as people’s cognition or their emotions) to make sense of behavior. The second direction suggests to look at the structure of environments to understand particular heuristics and cognitive biases that are domain-specific to those environments [57]. Our contribution relates to this second direction. We vary an individual’s social ecology (the specific stage game under consideration) and let a trait associated with rational learning evolve. Based on this setup, we explore how distinct environments shape individual adaptations.
@@
Our model is also directly related to two interesting studies by Szabó, Vukov, and Szolnoki [42, 43]. They also consider a model in which two traits co-evolve. These traits are the individuals’ strategies, and a noise parameter that is inversely related to our selection strength. Our work deviates from these studies in many ways. For example, we look at well-mixed populations while they study behavior on a regular lattice; we treat selection strength as a continuous variable while they consider a discrete set of possible noise parameters; we explore the entire range of $2\!\times\!2$ games whereas they consider a one-dimensional subspace of games. Most importantly, however, they consider simulations in which strategies and the noise parameter evolve at the same rate. Instead, we assume a separation of timescales. As a result, they do not address most of the questions that are central to our approach. For example, because they do not allow for lasting differences in an individual’s learning capabilities, they do not explore to which extent more effective learners are able to outperform their opponents. This question plays a crucial role in our setup, and it drives all of our subsequent results (such as the emergence of evolutionary branching, which they do not observe).
@@
Naturally, there are several ways how one could make our model more complex. For example, one could go beyond the symmetric one-shot games studied herein. Such work can shed light on the effects of asymmetric payoffs [33, 34, 35], repeated games [58], or heterogenous environments [59, 60]. In particular, such an approach could help us understand how people have developed psychological mechanisms to cope with the variability of situations they face. Another interesting direction would be to study the evolution of boundedly rational learning in structured rather than well-mixed populations. A framework of adaptive dynamics in such populations is readily available [61, 62]. This literature shows that network structure strongly affects the way individuals cooperate [63, 64, 65]. With a framework similar to ours, one could explore how network structure affects the way individuals learn. This work could investigate, for example, whether heterogeneous networks promote heterogeneity in people’s learning mechanisms. Finally, our results can also find future applications in multi-agent reinforcement learning. Specifically, one could conceive extensions where agents dynamically adapt their learning rates as a function of the dilemmas faced [66].
@@
We provide a detailed account of both the short-run learning dynamics and the long-run evolutionary dynamics in the Supporting Information. In the following we briefly sketch our methods for the long-run evolutionary dynamics. For that, we use a combination of individual-based simulations and a theoretical analysis in the infinite population limit.
@@
Our evolutionary simulations use a similar setup as Döbeli et al [67]. We start with a monomorphic population with trait $\beta=0$. At each time step, two individuals, $i$ and $j$, are randomly chosen from the population. These two individuals then compete for reproduction. To model this competition, each of the two individuals independently engage in a sample of three encounters with other random individuals. For each encounter, their payoff is computed with (3). By averaging over the three realizations, we compute each player’s average payoff $\pi_{i}$ and $\pi_{j}$. With probability $1/\big{(}1\!+\!\exp[-30(\pi_{i}\!-\!\pi_{j})]\big{)}$, individual $i$ then reproduces and $j$ dies. Otherwise, individual $j$ reproduces and $i$ dies. During reproduction, there is a small chance that an offspring inherits an imperfect copy of the parent trait (with a mutation rate of $\mu\!=\!0.005$). In that case, the mutation is uniformly distributed around the parent trait in an interval of size 0.1. We iterate this simulation step a large number of times. At regular time intervals, we record the composition of the population, and the average cooperation rates and payoffs. For the simulations in Fig. 3, we consider comparably large populations ($10^{4}$ in the prisoner’s dilemma and the snowdrift game; $2\!\cdot\!10^{4}$ in the stag-hunt game).
@@
For our analytical results, we use the adaptive dynamics formalism [38, 39, 40]. Here, we assume infinitely large populations and rare mutations of infinitesimal magnitude. Under these assumptions we can focus on mutant traits close-by to the resident, such that $y\!\approx\!x$. A key quantity to study such mutants is the local fitness gradient or selection gradient $D(x)$, defined by
@@
This condition determines whether the singular point can be reached at all, starting from a resident population sufficiently close-by. Moreover, the singular point is evolutionary stable if
@@
Data, Materials, and Software Availability. Simulations have been run with MATLAB. For our analytical results we used Mathematica. The respective simulation data and source code can be found in this OSF repository.

### v2

@@ Introduction
To study how the players’ payoff sensitivities $\beta$ might change over time, we use adaptive dynamics [27, 26, 34]. That is, we consider a large and well-mixed population of players (Fig. 1D). Each player randomly meets other population members to interact in a series of pairwise games. As a result, they obtain a payoff as described in the previous section on introspection dynamics. We define a player’s fitness to be its average payoff across all its interaction partners. Players with larger fitness are more likely to reproduce. The offspring of a reproducing player inherits the parent’s trait $\beta$ with probability $1\!-\!\mu$. With the complementary probability $\mu$, the offspring’s payoff sensitivity is slightly mutated by adding some small noise to the parent trait. Overall, these assumptions define a dynamical process on the population level, with $\beta$ being the evolving trait. This process is straightforward to simulate. We give a detailed description of the algorithm in the Material and Methods, and provide the respective code online.
@@
Initially, we assume that the population is monomorphic, with a payoff sensitivity of $\beta\!=\!0$. To gain some intuition of the further dynamics, we revisit the previous two examples. Fig. 3A shows simulations for the prisoner’s dilemma. Here, we observe that $\beta$ evolves towards ever-increasing values. This outcome is consistent with the naive expectation that less erratic learners should have a fitness advantage, and consistent with our analysis in Fig. 2A–D. For the other example, the snowdrift game, simulations are shown in Fig. 3C. Also here, larger payoff sensitivities are initially favored. After all, positive $\beta$ values ensure that less time is spent in off-equilibrium states (in fact, we prove in SI Section 1B.4 that for non-degenerate payoff matrices, $\beta$ always increases initially). On the other hand, the super-game depicted in Fig. 2H suggests that at some point, further increases in $\beta$ may no longer be favored, because it leads more sensitive players to spend more time in their less preferred equilibrium. Indeed, we observe the evolution of payoff sensitivities to stop at some value well below $\beta\!=\!6$. That is, in contrast to our naive expectation – but consistent with our analysis in Fig. 2E–H – we observe the evolution of finite $\beta$ values (or, of noisy learning).
@@
In this formula, $\Pi_{x}(y)$ is the payoff of a mutant with payoff sensitivity $\beta\!=\!y$ in a resident population with $\beta\!=\!x$. This payoff is given by Eq. [3]. The mutant can selectively invade the resident if and only if this invasion fitness is positive. In pairwise invasibility plots, the respective region for which $s_{x}(y)\!>\!0$ is highlighted in color. For the prisoner’s dilemma, this colored area comprises the entire space above the main diagonal (Fig. 3B). This means that independent of the exact payoff sensitivity $x$ of the resident, a mutant with trait $y$ can invade if and only if y>xy\!>\!x. This implies the evolution of ever-increasing payoff sensitivities.
@@
In contrast, in the snowdrift game in Fig. 3D, mutants with a higher payoff sensitivity are only favored when the resident’s sensitivity is sufficiently small (if $x\!<\!x^{*}$ for some threshold $x^{*}$). We can numerically compute this threshold by setting the so-called selection gradient equal to zero (see Material and Methods). For the depicted example, we obtain $x^{*}\!=\!5.58$, in line with the evolutionary simulations. The pairwise invasibility plot also indicates that this singular point is evolutionarily stable. Once the population settles at a payoff sensitivity of $\beta\!=\!x^{*}$, neither a more noisy nor a less noisy learner can invade (as indicated by the white neighborhood around the vertical dashed line).
@@
Interestingly, these two examples of an ever-increasing payoff sensitivity and an evolutionarily stable finite sensitivity are not the only possibilities. Instead, Fig. 3E shows a third scenario, using a stag-hunt game. Here, we observe evolutionary branching. At first, the monomorphic population again approaches a singular point at $x^{*}\!=\!3.01$. Once there, both mutant types can invade, those that are more and less noisy learners (indicated by the colored neighborhood of the vertical dashed line in Fig. 3F). This is the fingerprint of a branching point. From here, the population starts diverging in opposite directions. There is a sub-population for which sensitivities increase, and another one for which they decrease. The subsequent dynamics is more complex to describe analytically; we do so in the SI Section 1B.3 and Fig. S2. In the main text, we provide some intuition for why branching occurs in the first place. To this end, it is again instructive to look at the stationary distribution of introspection dynamics, for two different sensitivities $\beta$ (see Fig. S3A–D). When both players use the same $\beta$, a larger $\beta$ makes players choose the less efficient DD equilibrium more often (because defection is risk-dominant). If one player instead has a low $\beta$, both players occasionally coordinate on the more efficient CC equilibrium, while often avoiding harmful mis-coordination on the CD-outcome (see also SI Section 2C).
@@
Once the players’ sensitivity is allowed to evolve, its dynamics also has important implications on the population’s average payoff. For example in the prisoner’s dilemma, as individuals become more sensitive learners, they increasingly adopt the individually optimal strategy of defection. As a result, overall payoffs decrease in time (Fig. S4, left column). For the other two examples, the long-run effects on payoffs are more advantageous. Both for the snowdrift and the stag-hunt game, we observe an initially negative but eventually positive trend in the players’ payoffs (Fig. S4, middle and right column). In particular, the evolutionary branching in the stag-hunt game turns out to be both individually and collectively beneficial.
@@
So far, we have studied the evolution of payoff sensitivity for a few isolated $2\!\times\!2$ games. To identify more general patterns, in a next step, we systematically vary the parameters $A$ and $B$ of the stage game. For each resulting game, we compute the numerically exact position of the smallest singular point (if such a point exists) and its stability properties. The result is Fig. 4.
@@
Figure 4: Evolutionary dynamics across all social dilemmas. So far, we described three special instances of games; here we systematically repeat this analysis for all $2\!\times\!2$ games. For each pair of game parameters $A$ and $B$, we numerically check whether there exists a finite singular trait value $\beta^{*}$. If it exists, we explore whether the respective trait value is evolutionarily stable (orange) or a branching point (purple). The color gradient represents the position of the singular trait: the lighter, the larger the value of $\beta^{*}$. Regions without a finite singular trait are left white. Black symbols indicate the position of the three examples displayed in Fig. 3. Note that here we report the lowest singular point (the one which is reached when the process starts at $\beta=0$). However, there are games that permit two singular points. We show those in SI Section 2E.
Here, white regions reflect cases without a singular point. For the corresponding stage games, we predict sensitivities to increase indefinitely. Colored regions correspond to cases with at least one finite singular point. Here, we distinguish orange regions (where the smallest point is evolutionarily stable) and purple regions (where we predict branching). Based on this figure, we can make a number of observations.
@@
First, even though most prisoner’s dilemmas lead to ever-increasing payoff sensitivities (as in Fig. 3A), there are exceptions. These exceptions occur if the absolute value $|A|$ is small (if there is only a small temptation to defect when the co-player cooperates). These games permit a finite evolutionarily stable trait value of $\beta$. The underlying mechanism is different from the mechanism in the snowdrift game discussed earlier (Fig. 2E–H). Because $|A|$ is small compared to $B$, a co-player is more likely to occasionally switch to cooperation if the focal player cooperates. This positive influence on the co-player’s cooperation rate can make it worthwhile to have a finite $\beta$, and to cooperate occasionally (Fig. S3E–H).
@@
Second, in the snowdrift game, we observe evolution towards a finite payoff sensitivity for all games with B>AB\!>\!A. In addition, as we consider a wider range of parameter values than considered in Fig. 4, we find another region exhibiting evolution towards finite $\beta$, see Fig. S6. This second region is given by B<A1B\!<\!A\!-\!1; we provide a derivation of these conditions in the Supporting Information. In both regions, the mechanism favoring finite payoff sensitivities is the same: By reducing their sensitivity, learners ensure they spend more time in their individually preferred equilibrium.
@@
Third, in the stag-hunt game, we only observe evolutionary branching when it is risk-dominant to defect (that is, for A<BA\!<\!B). If instead cooperation is risk-dominant, the prevalence of CC increases as a function of the players’ payoff sensitivity. Mutual cooperation is the optimal outcome for both players; hence, an ever-increasing payoff sensitivity is individually and mutually beneficial.
Finally, in the harmony game quadrant, we observe a small region that permits the evolution of finite payoff sensitivities. The existence of this region is particularly puzzling. After all, in harmony games, cooperation is both individually and collectively optimal. We discuss one of these counterintuitive cases in Fig. S3I-L. For this case, the super-game for two given sensitivities can take the form of a prisoner’s dilemma: collectively, the two players reach better decisions when they both choose the larger $\beta$ value; yet individually, a player may deviate towards a smaller payoff sensitivity (as this incentivizes the more sensitive player to cooperate even more often). Interestingly, however, for each of these harmony games with a finite singular point, there exists in fact a second singular point. Hence, the adaptive dynamics for these harmony games is bistable. If players start out with small payoff sensitivities, they converge to a finite ESS. If their initial sensitivities are sufficiently large, however, their sensitivities further increase indefinitely (Fig. S5) – perhaps the more intuitive outcome in harmony games.
@@
Above, we documented several cases with an unexpected dynamics (e.g., evolution towards finite payoff sensitivities or evolutionary branching). In each case, the dynamics occurred because at least one player gained some strategic advantage from learning more erratically. In general, however, one might expect that these strategic advantages are strongest in pairwise games. Among two players, one player’s erratic behavior tends to have a comparably strong effect on how the co-player might react. This effect could wane in groups of size N>2N\!>\!2. To further explore the robustness of our results, in the Supporting Information we study the dynamics of three classes of public goods games. The three classes differ in the shape of the public good’s benefit function. Benefits are either linear, they show diminishing returns [29], or they take the form of a threshold function [6, 47]. While the computational steps involved in analyzing the adaptive dynamics are the same as before, the mathematical expressions now become more intricate. Importantly, however, the qualitative dynamics are similar to before. Depending on the payoff function, again we recover the three previously observed cases of indefinite evolution, convergence to a finite evolutionarily stable state, or evolutionary branching (Fig. S6, for N=3N\!=\!3).
Beyond introspection dynamics
In all our previous examples, we studied the evolution of finite payoff sensitivities under the assumption that individuals learn their strategies based on introspection dynamics. However, the learning literature knows of many other learning rules. To explore how our results generalize, in the SI Section 3, we numerically explore the adaptive dynamics of four other rules. All four rules can be regarded as special cases of experience-weighted attraction learning [13, 19, 48]: stochastic fictitious play, payoff-informed learning, and two different variants of reinforcement learning. In each case, we explore the evolution of the player’s motivation sensitivity $\lambda$ (which plays an analogous role to the parameter $\beta$ in introspection dynamics).
In general, for a given stage game, the four learning rules give rise to pairwise invasibility plots that can be rather different from each other, and different from introspection dynamics (Figs. S12-15). However, for each learning rule we find stage games for which $\lambda$ converges towards a finite value. Again, these cases are particularly common for stage games with more than one equilibrium (the snowdrift game and the stag-hunt game). In those cases, a finite $\lambda$ can again evolve as a means to settle at the player’s more preferred equilibrium (Figs. S9-S11). In our view, the mere existence of these finite singular points (across all considered learning rules) is remarkable. It suggests a surprisingly pervasive strategic advantage of noisy learning in certain types of stage games.
@@ Discussion
Evolutionary game theory explores how strategies and traits change over time [45, 12, 60]. Traditionally, such an evolutionary approach can be motivated in two ways. On the one hand, the respective dynamical equations may describe long-run biological evolution. Here, individuals with better strategies are assumed to reproduce more often [39, 33]. On the other hand, evolutionary models can also be interpreted as a representation of an individual’s short-run learning process. Such models assume that better strategies are more readily adopted [9, 55, 59]. Herein, we blend these two interpretations. We ask how a long-term biological process may shape the effectiveness of individual learning. To this end, we study the evolution of a key quantity in stochastic learning models, the players’ payoff sensitivity (which is often referred to as ‘selection strength’ in the related literature). This quantity determines to which extent payoff advantages sway individuals to change their strategies during the learning process. When individuals are comparably insensitive, they tend to ignore their strategies’ payoffs [44, 58, 41, 3, 54]. Here, strategy choices are mostly determined by chance. As payoff sensitivities become larger, individuals become increasingly swayed by (even minor) payoff differences.
@@
Motivated by these observations, herein we interpret payoff sensitivity as an individual trait that is related to the amount of noise in an individual’s updating process. We use this perspective to study the evolution of noisy learning. To this end, we consider a model with two time scales. In the short run, individuals have fixed but possibly different payoff sensitivities. We explore the resulting learning processes with introspection dynamics [15, 16, 32, 40, 63, 36, 52, 50]; for the $2\!\times\!2$ games we focus on, this process is equivalent to the logit-response dynamics studied in economics [16, 5]. We use this short-term perspective to ask whether individuals who learn less erratically are able to take advantage of their opponents.
@@
In the long run, we also allow payoff sensitivities to evolve. Here, naively one might expect ever-increasing payoff sensitivities. After all, large payoff sensitivities allow individuals to better distinguish profitable strategies from unprofitable ones. Surprisingly, however, we identify several important game classes in which payoff sensitivity converges towards a finite value. When learning more erratically, individuals occasionally use strategies that run counter to their immediate interests. In turn, players who are more sensitive to payoffs are swayed to react and to coordinate on an equilibrium that eventually benefits the noisier learner. These findings have interesting parallels in the biology literature on the red-king effect [8, 17, 61]. That effect suggests that in mutualisms among two co-evolving species, it can be the slower-evolving species that eventually has an evolutionary advantage.
@@
Our model is also directly related to two interesting studies by Szabó, Vukov, and Szolnoki [56, 57]. They also consider a model in which two traits co-evolve. These traits are the individuals’ strategies and a noise parameter that is inversely related to our payoff sensitivity. Our work deviates from these studies in many ways. For example, we look at well-mixed populations while they study behavior on a regular lattice; we treat payoff sensitivity as a continuous variable while they consider a discrete set of possible noise parameters; we explore the entire range of $2\!\times\!2$ games, whereas they consider a one-dimensional subspace of games. Most importantly, however, they consider simulations in which strategies and the noise parameter evolve at the same rate. Instead, we assume a separation of timescales. As a result, they do not address most of the questions that are central to our approach. For example, because they do not allow for lasting differences in an individual’s learning capabilities, they do not explore to which extent more sensitive learners are able to outperform their opponents. This question plays a crucial role in our setup, and it drives all of our subsequent results (such as the emergence of evolutionary branching, which they do not observe).
We have already sketched possible generalizations of our baseline model. In particular, in the Supporting Information we provide more details on how the model can be extended to cover multiplayer interactions (SI Section 2F) or alternative learning rules (SI Section 3). Several other extensions are possible too. For example, herein, we considered individuals who engage in a series of one-shot (non-repeated) games. In SI Section 2G, we sketch instead how to capture strategies of direct reciprocity [25] within our setup. To this end, we consider the pairwise dynamics when individuals choose between ALLD and Tit-for-Tat. Surprisingly, we find that depending on how individuals discount the future, already these two strategies can generate all of our previously described qualitative scenarios (see Fig. S7). Future work could provide a more comprehensive analysis that also includes other repeated-game strategies. Other more complex setups that one could study within our framework are asymmetric games [32, 63, 36], and heterogeneous environments [7, 14]. Such an approach could help us understand how people have developed psychological mechanisms to cope with the variability of situations they face.
@@
Yet another interesting direction would be to study the evolution of noisy learning in structured rather than well-mixed populations. A framework of adaptive dynamics in such populations is readily available [4, 31]. The literature shows that network structure strongly affects the way individuals cooperate [46, 51, 49]. With a framework similar to ours, one could explore how network structure affects the way individuals learn. This work could investigate, for example, whether heterogeneous networks promote heterogeneity in people’s learning mechanisms. Finally, our results can also find future applications in multi-agent reinforcement learning. Specifically, one could conceive extensions where agents dynamically adapt their learning rates as a function of the dilemmas faced [10].
@@
We provide a detailed account of both the short-run learning dynamics and the long-run evolutionary dynamics in the Supporting Information. In the following, we briefly sketch our methods for the long-run evolutionary dynamics. For that, we use a combination of individual-based simulations and a theoretical analysis in the infinite population limit.
@@
Our evolutionary simulations use a similar setup as Döbeli et al [18]. We start with a monomorphic population with trait $\beta=0$. At each time step, two individuals, $i$ and $j$, are randomly chosen from the population. These two individuals then compete for reproduction. To model this competition, each of the two individuals independently engages in a sample of three encounters with other random individuals. For each encounter, their payoff is computed with (3). By averaging over the three realizations, we compute each player’s average payoff $\pi_{i}$ and $\pi_{j}$. With probability $1/\big(1\!+\!\exp[-30(\pi_{i}\!-\!\pi_{j})]\big)$, individual $i$ then reproduces and $j$ dies. Otherwise, individual $j$ reproduces and $i$ dies. During reproduction, there is a small chance that an offspring inherits an imperfect copy of the parent trait (with a mutation rate of $\mu\!=\!0.005$). In that case, the mutation is uniformly distributed around the parent trait in an interval of size 0.1. We iterate this simulation step a large number of times. At regular time intervals, we record the composition of the population and the average cooperation rates and payoffs. For the simulations in Fig. 3, we consider comparably large populations ($10^{4}$ in the prisoner’s dilemma and the snowdrift game; $2\!\cdot\!10^{4}$ in the stag-hunt game).
@@
For our analytical results, we use the adaptive dynamics formalism [27, 26, 34]. Here, we assume infinitely large populations and rare mutations of infinitesimal magnitude. Under these assumptions, we can focus on mutant traits close to the resident, such that $y\!\approx\!x$. A key quantity to study such mutants is the local fitness gradient or selection gradient $D(x)$, defined by
@@
This condition determines whether the singular point can be reached at all, starting from a resident population sufficiently close by. Moreover, the singular point is evolutionary stable if
@@
Data, Materials, and Software Availability. Simulations have been run with MATLAB. For our analytical results, we used Mathematica. The respective simulation data and source code can be found in this OSF repository.